Phosphoric acid can be prepared from phosphorus triiodide according to the reaction:

    Pl3 (s)  +  H2O(l) -----------> H3PO3(aq) + HI(g)

If 150 g of PI3 (MM = 411.7 g/mol) is added to 250 mL of H2O (g/mol), d = 1.00 g/mL), Identify which will be limiting and excess reagents. 



1
Expert's answer
2021-11-22T17:17:03-0500

PI3 + 3H2O → H3PO3 + 3HI

n(PI) "= \frac{150}{411.7} = 0.364 \;mol"

m(H2O) "= 1.00 \times 250 = 250 \;g"

M(H2O) = 18.0 g/mol

n(H2O) "= \frac{250}{18.0} = 13.88 \;mol"

According to the reaction for one mole of PI we need 3 moles of water. So, PI is a limiting reactant and water is an excess reactant.